mikewwe13:

What are the x-intercepts of the quadratic function? f(x) = x^2 − 3x − 10 Enter your answers in the boxes. ______ and ______

mikewwe13:

@Vocaloid

Vocaloid:

start by factoring it

mikewwe13:

(5, 0) and (-2, 0)

Vocaloid:

yeah that's right if the problem doesn't accept that as your answer, it might want just "5" and "-2" instead, I'm not sure

mikewwe13:

you think ?

Vocaloid:

try entering (5, 0) and (-2, 0) first, if it doesn't accept that, enter 5 and -2

Vocaloid:

your math is correct

mikewwe13:

Let ​f(x) = x^2 + 14x + 48​. What are the zeros of the function? Enter your answers in the boxes. ______ and _______

Vocaloid:

same thing as before, start factoring

mikewwe13:

f(x) = (x - 4) (x + 6)

Vocaloid:

that's not quite right, check your factoring again hint - pick two integers that multiply together to get 48 and add to get 14

mikewwe13:

12 x 4

Vocaloid:

not quite 12 + 4 is not 14

mikewwe13:

the answers are -6 and 4

Vocaloid:

nope -6 + 4 is not 14 remember the sum must be ***14***

mikewwe13:

no that's x intercepts

Vocaloid:

does (x+6)(x-4) = x^2 + 14x + 48​? if not, then 6 and -4 are NOT the intercepts.

mikewwe13:

(4, 0) and (-6, 0)

mikewwe13:

and yes

Vocaloid:

use foil to expand (x+6)(x-4) you will see that it is not equal to x^2 + 14x + 48

Vocaloid:

again, you must pick two numbers that *** add up to 14 *** and *** multiply to get 48 ***

mikewwe13:

2x for 14 and 24 for 48

Vocaloid:

nope, keep trying, we are looking for two integers (not variables)

Vocaloid:

start by listing pairs of numbers that multiply to get 48

mikewwe13:

6 x 8 = 48, 6 and 8 add up to 14 and multiply to 48

Vocaloid:

good so we factor to get (x+6)(x+8) = 0 what are the x-intercepts?

mikewwe13:

x = -6, -8

Vocaloid:

good, so (-6,0) and (-8,0) are your intercepts

mikewwe13:

What are the x-intercepts of the quadratic function? f(x) = x^2 + 2x − 24 Enter your answers in the boxes. ____ and _____

mikewwe13:

did we did this already ?

Vocaloid:

not yet try picking two numbers that multiply to get -24 and add to get 2

mikewwe13:

the answers is -6 and 4

Vocaloid:

for the intercepts? yeah, (-6,0) and (4,0)

mikewwe13:

thank you so very much